In the Old West, settlers often fashioned tents out of a piece of cloth thrown over tent poles and then secured to the ground with stakes forming an isosceles triangle. How long would the cloth have to so that the opening of the tent was 4 meters high and 3 meters wide?
I tryed this problem thinking that you could solve using the Pythagorean Theorm but you can only use that if it is a right triangle.

Answer by stanbon(75887)

isosceles triangle. How long would the cloth have to so that the opening of the tent was 4 meters high and 3 meters wide?
Draw the picture and imagine the open door is being
held up by a pole.
The pole is the altitute of the isosceles triangle.
Do you see the two fight angles formed at the base of
the pole?
Now do you see the two right triangles that are formed?
Use Pythagoras to find 1/2 the length of the cloth, as follows:
cloth^2 = base^2 + altitude^2
c^2 = (3/2)^2 + 4^2
c^2 = 9/4 +64/4= 73/4
c= (1/2)sqrt(73) (that's the length of 1/2 the cloth
2c= sqrt(73) meters (that's the length of the whole cloth
Cheers,
Stan H>